ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 0, 3 Sahab, Salem, King Abdul-Aziz University, Saudi-Arabia Mathematics basic skills for grades 1 to 3 As a result of a national research project in Saudi Arabia, a single list of the basic mathematics skills has been established for each of the first three elementary grades. Stnadardized tests were made for each of these grades, and shall be applied to a random sample of pupils sometimes in May. The results of these tests will be discussed and analyzed to see how far is the gernal population of pupils are learning these important basic skills. Of course, most of these skill deal with arithematics. But surely, lots of old and recent research has proven that later mathematics learning and understanding depend largely on what happens in these early years of every child.